# How Muon’s Spectral Design Benefits Generalization: A Study on Imbalanced Data

### v1

@@ 1 Introduction
Spectrum-aware optimizers such as Shampoo (Gupta et al.,, 2018), Muon (Pethick et al.,, 2025; Jordan et al.,, 2024) have recently gained significant traction in the deep learning community, delivering substantial training speedups for deep classifiers (Jordan et al.,, 2024) and transformer language models (Vyas et al.,, 2025; Liu et al.,, 2025) compared to standard methods like SGD (Robbins and Monro,, 1951) with momentum or even Adam (Kingma and Ba,, 2014). The key distinction lies in how these methods treat neural network parameters: while SGD and Adam operate entry-wise on vectorized parameters, Shampoo and Muon work directly with matrix-valued parameters (e.g., weight matrices and attention matrices) at the layer level (Carlson et al.,, 2015; Large et al.,, 2024). This matrix-level approach intuitively enables optimization trajectories that entry-wise methods cannot achieve. Despite their empirical success, a fundamental question remains unanswered: when do spectrum-aware optimizers generalize better than standard methods?
The challenge is substantial. Even well-established optimizers like Adam, despite more than a decade of practical dominance, remain less well understood than SGD, whose Euclidean GD trajectory is well-characterized both statistically and algorithmically. Recent theoretical progress has begun to illuminate these methods through the lens of implicit bias. For instance, while the implicit bias of GD toward $\ell_{2}$-norm max-margin classifiers is well-established (Soudry et al.,, 2018; Ji and Telgarsky,, 2018), recent work proved that Adam converges toward $\ell_{\infty}$-norm max-margin classifiers in linear settings (Zhang et al.,, 2024; Xie and Li,, 2024). Conceptually, this difference can be understood by noting that Adam reduces to SignGD when momentum and preconditioning histories are ignored (i.e., $\beta$ parameters set to zero). This reduction has often been leveraged before to study Adam’s properties, since SignGD is simpler to analyze, e.g., Kunstner et al., (2024); Vasudeva et al., 2025b .
A similar reduction exists for spectrum-aware optimizers: Shampoo without preconditioning history and Muon with exact matrix operations reduce to Spectral Gradient Descent (SpecGD)(Pethick et al.,, 2025; Bernstein and Newhouse,, 2024), where each update step is ${\bm{U}}{\bm{V}}^{T}$, with ${\bm{U}}\bm{\Sigma}{\bm{V}}^{T}$ denoting the truncated SVD of the gradient. Thus, just as SignGD serves as the canonical form for understanding Adam, SpecGD serves as a canonical form for understanding Shampoo and Muon. Moreover, both SignGD and SpecGD are instances of normalized steepest descent with respect to $\ell_{\infty}$ and spectral norms, respectively (Bernstein,, 2009; Carlson et al.,, 2015; Pethick et al.,, 2025).
@@
Recently, Fan et al., (2025); Tsilivis et al., (2024) characterized SpecGD’s optimization bias in linear multi-class and homogeneous binary classification, showing that it drives weights toward a max-margin classifier with respect to the spectral norm. However, these results have two limitations. First, they only describe the algorithm’s behavior in the terminal phase of training, which may not reflect practical deep learning scenarios that often employ early stopping. Second, and more importantly, implicit bias results provide no direct guarantees about generalization performance—the ultimate objective in machine learning. For instance, even in linear settings, the minimum spectral norm solution may not be unique. Recent work has also studied Muon and Shampoo, but primarily from the perspective of optimization properties (Pethick et al.,, 2025; Morwani et al.,, 2024; Vyas et al.,, 2025; Li and Hong,, 2025; Chang et al.,, 2025; Chen et al.,, 2025) or scalability (Liu et al.,, 2025; Boreiko et al.,, 2025), leaving their generalization behavior and underlying mechanisms comparatively underexplored.
@@
Imbalanced Data as Playground. We introduce imbalanced data as a testbed for studying SpecGD’s potential generalization advantages. Fig.˜1 provides a concrete demonstration: we train an MLP on Colored-MNIST (Arjovsky et al.,, 2020; Pezeshki et al.,, 2021) under severe group imbalance, where each digit appears in its majority color $99\%$ of the time during training, creating a strong spurious digit-color correlation (see Sec.˜D.1 for details). At test time, we evaluate on majority and minority groups with same and opposite digit-color associations, respectively. The results are revealing: whereas Muon improves on both groups early in training, its Frobenius and $\max$ norm counterparts, namely normalized momentum descent (NMD) and Signum (Bernstein et al.,, 2018), respectively, take much longer to improve on the minority group. We observe a similar effect on the CIFAR dataset (Krizhevsky et al.,, 2009) with class imbalance (see Sec.˜2 for details).
@@ 2 Results on CIFAR with STEP-Imbalance
In this section, we conduct a preliminary experiment to investigate whether Muon’s superiority observed in the group-imbalanced setting in Fig.˜1 also persists under class imbalance. We consider step-imbalanced CIFAR-10 and CIFAR-100 (Krizhevsky et al.,, 2009) with imbalance ratio $20$ (majorities contain $20\times$ more samples than minorities). We train ResNet-18/50 on CIFAR-10/100 with three optimizers: NMD, Signum, and Muon (see Sec.˜D.2 for details and additional results with lower imbalance ratios). As shown in Fig.˜3, Muon achieves superior minority-class performance compared to NMD and Signum. While all optimizers achieve similar majority-class accuracy, Muon more effectively reduces the performance gap between minority and majority classes, particularly early on in training. This result further suggests that spectral methods may offer advantages in imbalanced settings, motivating our subsequent theoretical investigation into the mechanisms underlying this behavior.
@@ 3 Linear Model on Class-Imbalanced Data
We denote matrices, vectors and scalars by $\bm{A}$, $\bm{a}$, and $a$, respectively. We denote the $(i,j)$-th entry of matrix $\bm{A}$ as $A[i,j]$. Let $\lVert\cdot\rVert_{F}$, $\lVert\cdot\rVert_{2}$, and $\lVert\cdot\rVert_{\max}$ denote the Frobenius, spectral and $\max$ norms, respectively, where $\lVert\bm{A}\rVert_{\max}:=\max_{i,j}A[i,j]$. Let $\lVert\bm{a}\rVert_{2}$ denote the $\ell_{2}$ norm of $\bm{a}$. $\mathds{1}\left[\cdot\right]$ denotes the indicator function, e.g., $\mathds{1}\left[a\geq b\right]=1$ if $a\geq b$ and 0 otherwise.
@@
As discussed in Sec.˜1, NGD, SignGD and SpecGD are instances of normalized steepest descent Eq.˜1 with respect to Frobenius, $\max$ and spectral norm, respectively (also see App.˜A).
@@ 3.1 Experimental Results
For the experiments, we consider a heavy-tailed class-imbalanced setting by choosing $p_{c}\!\propto\!\tfrac{1}{c}$, and we sample each $\bm{\mu}_{c}$ independently from a zero-mean isotropic Gaussian distribution and normalize it. We use $20$ classes, $100$ samples, $d=200$ and $\sigma_{x}=0.1$. We initialize $\bm{W}_{0}$ by sampling each entry independently from $\mathcal{N}(0,\tfrac{1}{d})$. We use learning rates $0.025$, $0.005$ and $5\times 10^{-4}$ for NGD, SignGD and SpecGD, respectively. These choices are made such that train loss curves of the algorithms are comparable. The training is stopped when the norm of the gradient for SpecGD falls below $10^{-6}$, as small gradients can introduce numerical imprecision in the SVD computation.
Fig.˜4 shows results of training a linear model in this setting using NGD/SignGD/SpecGD to minimize cross-entropy loss. We observe that for all three update rules, as train loss approaches 0, the model converges to a solution that maximizes the margin defined with respect to the corresponding norm (Fan et al.,, 2025). However, comparing test performance, early-stopped SpecGD attains higher class-balanced and worst-class test accuracies compared to NGD or SignGD at any stopping point.
@@ 3.2 Theoretical Analysis
Under this condition applied to the empirical moment matrices, Saxe et al., (2013); Gidel et al., (2019) derive closed-form training dynamics of two-layer linear networks. Here, we instead apply this condition on population statistics making it possible to study test statistics. In the lemma below we show that our data model (DM) satisfies this condition. See Sec.˜B.1 for the proof.
@@
The population moment matrices of data model (DM) satisfy Condition 1, with $s^{\texttt{yx}}_{c}=p_{c}$ for $c\in[k]$, $s^{\texttt{xx}}_{c}=\mu^{2}p_{c}+\sigma_{x}^{2}$ for $1\leq c\leq k$, and $s^{\texttt{xx}}_{c}=\sigma_{x}^{2}$ for $k<c\leq d$.
@@
Under Condition 1, Saxe et al., (2013) (see also Sec.˜B.2) shows that when initialized at zero and run with sufficiently small step size, GD iterates $\bm{W}_{t}$ are such that $\overline{\bm{W}}_{t}$ is diagonal at each iteration with diagonal entries evolving as (and approximation becoming accurate for gradient flow (GF) $\eta\!\rightarrow\!0$):
@@
Assume zero initialization $\bm{W}_{\bm{0}}={\bm{0}}$ and Condition 1 holds. Then, for SpecGD, at each iteration, $\bm{W}_{t}={\bm{U}}\overline{\bm{W}}_{t}{\bm{V}}^{\top}$ where $\overline{\bm{W}}_{t}$ is zero except its main diagonal along which the entries evolve as follows for $c\in[k]$:
@@
Comparing the above two displays, which contrast GD’s and SpecGD’s iterate evolution, shows the following. Although both methods asymptotically converge to the same solution, their trajectories differ significantly. GD learns component $c$ at a rate proportional to $s^{\texttt{xx}}_{c}$, meaning more dominant spectral components are learned faster. In contrast, SpecGD learns all components at the same rate until each individual value saturates and converges to its terminal value. This difference is also demonstrated in Fig.˜2 (top), which tracks $\overline{W}_{t}[c,c]$ for GD and SpecGD in a class-imbalanced setting (data model (DM)) with $\eta=0.01$, $d=k=3$, $\mu=1$, $\sigma_{x}^{2}=0.125$, and $p_{1},p_{2},p_{3}$ set as $0.5,0.3,0.2$, respectively. In App.˜C, we compare these dynamics of GD and SpecGD derived under Condition 1 to a more general setting with the same data model but with random initialization and finite samples, and show that the theoretically derived dynamics closely match those observed empirically.
@@
We now show that this property of SpecGD to learn concepts at equal rate translates to superior generalization compared to GD in an imbalanced setting where least-significant spectral components of the moment matrices are associated to minority classes. Concretely, under data model (DM), the $k$ first eigenvectors of $\bm{\Sigma}_{{\bm{x}}{\bm{x}}}$ align with the class-mean vectors, ordered in decreasing class prior probability (see proof of Lemma 3.1). Intuitively, then, learning spectral components earlier during training should translate to generalization gains. Indeed, in Fig.˜2 (bottom), we observe that SpecGD attains lower minority-class and class-balanced loss compared to GD. The following theorem formalizes this intuition. We use continuous-time versions of GD and SpecGD here, i.e., gradient flow (GF) and spectral GF (SpecGF) [^2] The proof also works for discrete time versions, but the dynamics are more precise for gradient flow.. The dynamics for GF are characterized by the approximation in Eq.˜2, while the specGF dynamics follow Proposition˜1, both with $\eta=1$. We use $\bm{W}(t)$ and $\overline{\bm{W}}(t)$ to denote the continuous-time iterates and the corresponding singular value matrix, respectively.
@@
Next, we investigate whether these gains of SpecGD arise because it uses normalized updates while GD does not. To account for this difference, we compare SpecGD with NGD, the normalized version of GD. We first simulate the dynamics of NGD in the setting of Fig.˜2, and find that although NGD converges much faster than GD, its iterates $\overline{\bm{W}}_{t}$ follow the same trajectory as GD (Fig.˜2 (top)). Importantly, we find that SpecGD still outperforms NGD in terms of both minority-class and class-balanced loss (Fig.˜2 (bottom)). We formalize and prove this observation in the theorem below. See Sec.˜B.4 for the complete version and the proof.

### v2

@@ 1 Introduction
Spectrum-aware optimizers such as Shampoo (gupta2018shampoo), Muon (pethick2025trainingdeeplearningmodels; jordan2024muon) have recently gained significant traction in the deep learning community, delivering substantial training speedups for deep classifiers (jordan2024muon) and transformer language models (vyas2025soap; liu2025muon) compared to standard methods like SGD (robbins1951stochastic) with momentum or even Adam (kingma2014adam). The key distinction lies in how these methods treat neural network parameters: while SGD and Adam operate entry-wise on vectorized parameters, Shampoo and Muon work directly with matrix-valued parameters (e.g., weight matrices and attention matrices) at the layer level (Carlson2015PreconditionedSD; large2024scalable). This matrix-level approach intuitively enables optimization trajectories that entry-wise methods cannot achieve. Despite their empirical success, a fundamental question remains unanswered: when do spectrum-aware optimizers generalize better than standard methods?
The challenge is substantial. Even well-established optimizers like Adam, despite more than a decade of practical dominance, remain less well understood than SGD, whose Euclidean GD trajectory is well-characterized both statistically and algorithmically. Recent theoretical progress has begun to illuminate these methods through the lens of implicit bias. For instance, while the implicit bias of GD toward $\ell_{2}$-norm max-margin classifiers is well-established (soudry2018implicit; ji2018gradient), recent work proved that Adam converges toward $\ell_{\infty}$-norm max-margin classifiers in linear settings (adam2024implicit1; xie2024implicit). Conceptually, this difference can be understood by noting that Adam reduces to SignGD when momentum and preconditioning histories are ignored (i.e., $\beta$ parameters set to zero). This reduction has often been leveraged before to study Adam’s properties, since SignGD is simpler to analyze, e.g., kunstner2024heavy; vasudeva2025richsimpleimplicitbias.
A similar reduction exists for spectrum-aware optimizers: Shampoo and Muon without accumulations and with exact matrix operations reduce to Spectral Gradient Descent (SpecGD) (pethick2025trainingdeeplearningmodels; bernstein2024old), where each update step is ${\bm{U}}{\bm{V}}^{T}$, with ${\bm{U}}\bm{\Sigma}{\bm{V}}^{T}$ denoting the truncated SVD of the gradient (see Sec.˜3 for details). Thus, just as SignGD serves as the canonical form for understanding Adam, SpecGD serves as a canonical form for understanding Shampoo and Muon. Moreover, both SignGD and SpecGD are instances of normalized steepest descent with respect to $\max$ and spectral norms, respectively (bernstein2009matrix; Carlson2015PreconditionedSD; pethick2025trainingdeeplearningmodels).
@@
Recently, Chen_NSD; tsilivis2024flavors characterized SpecGD’s optimization bias in linear multi-class and homogeneous binary classification, showing that it drives weights toward a max-margin classifier with respect to the spectral norm. However, these results have two limitations. First, they only describe the algorithm’s behavior in the terminal phase of training, which may not reflect practical deep learning scenarios that often employ early stopping. Second, and more importantly, implicit bias results provide no direct guarantees about generalization performance—the ultimate objective in machine learning. For instance, even in linear settings, the minimum spectral norm solution may not be unique. Recent work has also studied Muon and Shampoo, but primarily from the perspective of optimization properties (pethick2025trainingdeeplearningmodels; morwani2024newperspectiveshampoospreconditioner; vyas2025soap; li2025note; chang2025convergencemuon; chen2025muonoptimizesspectralnorm) or scalability (liu2025muon; boreiko2025towards), leaving their generalization behavior and underlying mechanisms comparatively underexplored.
@@
Imbalanced Data as Playground. We introduce imbalanced data as a testbed for studying SpecGD’s potential generalization advantages. Fig.˜1 provides a concrete demonstration: we train an MLP on Colored-MNIST (irm; gs) under severe group imbalance, where each digit appears in its majority color $99\%$ of the time during training, creating a strong spurious digit-color correlation (see Sec.˜D.1 for details). At test time, we evaluate on majority and minority groups with same and opposite digit-color associations, respectively. The results are revealing: whereas Muon improves on both groups early in training, its Frobenius and $\max$ norm counterparts, namely normalized momentum descent (NMD) and Signum (bernstein2018signsgd), respectively, take much longer to improve on the minority group. We observe a similar effect on the CIFAR dataset (krizhevsky2009learning) with class imbalance (see Sec.˜2 for details).
@@ 2 Results on CIFAR with STEP-Imbalance
In this section, we conduct a preliminary experiment to investigate whether Muon’s superiority observed in the group-imbalanced setting in Fig.˜1 also persists under class imbalance. We consider step-imbalanced CIFAR-10 and CIFAR-100 (krizhevsky2009learning) with imbalance ratio $20$ (majorities contain $20\times$ more samples than minorities). We train ResNet-18/50 on CIFAR-10/100 with three optimizers: NMD, Signum, and Muon (see Sec.˜D.2 for details and additional results with lower imbalance ratios). As shown in Fig.˜3, Muon achieves superior minority-class performance compared to NMD and Signum. While all optimizers achieve similar majority-class accuracy, Muon more effectively reduces the performance gap between minority and majority classes, particularly early on in training. This result further suggests that spectral methods may offer advantages in imbalanced settings, motivating our subsequent theoretical investigation into the mechanisms underlying this behavior.
@@ 3 Linear Model on Class-Imbalanced Data
We denote matrices, vectors and scalars by $\bm{A}$, $\bm{a}$, and $a$, respectively. We denote the $(i,j)$-th entry of matrix $\bm{A}$ as $A[i,j]$. Let $\lVert\cdot\rVert_{F}$, $\lVert\cdot\rVert_{\max}$, and $\lVert\cdot\rVert_{2}$ denote the Frobenius, $\max$ and spectral norms, respectively, where $\lVert\bm{A}\rVert_{\max}:=\max_{i,j}|A[i,j]|$. Let $\lVert\bm{a}\rVert_{2}$ denote the $\ell_{2}$ norm of $\bm{a}$. $\mathds{1}\left[\cdot\right]$ denotes the indicator function, e.g., $\mathds{1}\left[a\geq b\right]=1$ if $a\geq b$ and 0 otherwise. For matrices $\bm{A},\bm{B}$, the inner product $\left\langle\bm{A},\bm{B}\right\rangle=\operatorname{tr}(\bm{A}\bm{B}^{\top})$.
@@
[…]
$\lVert\cdot\rVert_{2}$ SpecGD: $\mathbf{\Delta}_{t}={\bm{U}}_{t}{\bm{V}}_{t}^{\top}$ Muon: $\mathbf{\Delta}_{t}=\widetilde{{\bm{U}}}_{t}\widetilde{{\bm{V}}}_{t}^{\top}$
@@ 3.1 Experimental Results
For the experiments, we consider a heavy-tailed class-imbalanced setting by choosing $p_{c}\!\propto\!\tfrac{1}{c}$, and we sample each $\bm{\mu}_{c}$ independently from a zero-mean isotropic Gaussian distribution and normalize it. We use $20$ classes, $100$ samples, $d=200$ and $\sigma_{x}=0.1$. We initialize the weight matrix $\bm{W}_{0}$ by sampling each entry independently from $\mathcal{N}(0,\tfrac{1}{d})$. We use learning rates $0.025$, $0.005$ and $5\times 10^{-4}$ for NGD, SignGD and SpecGD, respectively. These choices are made such that train loss curves of the algorithms are comparable. The training is stopped when the norm of the gradient for SpecGD falls below $10^{-6}$, as small gradients can introduce numerical imprecision in the SVD computation.
Fig.˜4 shows results of training a linear model in this setting using NGD/SignGD/SpecGD to minimize cross-entropy loss. We observe that for all three update rules, as train loss approaches 0, the iterates $\bm{W}_{t}$ converge to a solution that maximizes the margin defined with respect to the corresponding norm (Chen_NSD). However, comparing test performance, early-stopped SpecGD attains higher class-balanced and worst-class test accuracies compared to NGD or SignGD at any stopping point.
@@ 3.2 Theoretical Analysis
Under this condition applied to the empirical moment matrices, saxe2013exact; gidel2019implicit derive closed-form training dynamics of two-layer linear networks. Here, we instead apply this condition on population statistics making it possible to study test statistics. In the lemma below, we show that our data model (DM) satisfies this condition. See Sec.˜B.1 for the proof.
@@
The population moment matrices of data model (DM) satisfy Condition 1, with $s^{\texttt{yx}}_{c}=\mu p_{c}$ for $c\in[k]$, $s^{\texttt{xx}}_{c}=\mu^{2}p_{c}+\sigma_{x}^{2}$ for $1\leq c\leq k$, and $s^{\texttt{xx}}_{c}=\sigma_{x}^{2}$ for $k<c\leq d$.
This result shows that the data model (DM) is one such setting where Condition 1 strictly holds. Notably, in gidel2019implicit, the authors validate whether a weaker version of this condition is satisfied for some datasets used in practice. Specifically, they check whether there exist orthonormal matrices ${\bm{U}},{\bm{V}}$ such that the empirical moment matrices are jointly diagonalizable as ${\bm{U}}\bm{\Sigma}_{{\bm{y}}{\bm{x}}}{\bm{V}}^{\top}$ and ${\bm{V}}(\bm{\Sigma}_{{\bm{x}}{\bm{x}}}+\bm{B}){\bm{V}}^{\top}$, respectively, where $\lVert\bm{B}\rVert$ is much smaller than $\lVert\bm{\Sigma}_{{\bm{x}}{\bm{x}}}\rVert$. They find that on datasets like MNIST, CIFAR, etc., the ratio is indeed small (see Table 1 in gidel2019implicit), so this can be considered as a reasonable condition. We consider Condition 1, where $\lVert\bm{B}\rVert=0$, mainly to make the analysis more tractable. See App.˜C for further discussion, and experimental results that examine the effect of relaxing this condition.
@@
Under Condition 1, saxe2013exact (see also Sec.˜B.2) shows that when initialized at zero and run with sufficiently small step size $\eta<\min_{c\in[k]}(s^{\texttt{xx}}_{c})^{-1}$, GD iterates $\bm{W}_{t}$ are such that $\overline{\bm{W}}_{t}$ is diagonal at each iteration with diagonal entries evolving as (and approximation becoming accurate for gradient flow (GF) $\eta\!\rightarrow\!0$):
@@
Assume zero initialization $\bm{W}_{\bm{0}}={\bm{0}}$ and Condition 1 holds. Then, for SpecGD with step size $\eta<\min_{c\in[k]}\tfrac{s^{\texttt{yx}}_{c}}{s^{\texttt{xx}}_{c}}$, at each iteration, $\bm{W}_{t}={\bm{U}}\overline{\bm{W}}_{t}{\bm{V}}^{\top}$ where $\overline{\bm{W}}_{t}$ is zero except its main diagonal along which the entries evolve as follows for $c\in[k]$:
@@
Comparing the above two displays, which contrast GD’s and SpecGD’s iterate evolution, shows the following. Although both methods asymptotically converge to the same solution, their trajectories differ significantly. GD learns component $c$ at a rate proportional to $s^{\texttt{xx}}_{c}$, meaning more dominant spectral components are learned faster. In contrast, SpecGD learns all components at the same rate until each individual value saturates and converges to its terminal value. This difference is also demonstrated in Fig.˜2 (top), which tracks $\overline{W}_{t}[c,c]$ for GD and SpecGD in a class-imbalanced setting (data model (DM)) with $\eta=0.01$, $d=k=3$, $\mu=1$, $\sigma_{x}^{2}=0.125$, and $p_{1},p_{2},p_{3}$ set as $0.5,0.3,0.2$, respectively.
In App.˜C, we compare the dynamics of GD and SpecGD derived under Condition 1 to a more general setting using the same data model but with finite samples and random initialization, and show that the theoretically derived dynamics closely match those observed empirically (see Fig.˜11). Additionally, in App.˜C, we also compare the dynamics of Muon (with $\beta=0.9$) to those of SpecGD, both when Condition 1 holds (see Fig.˜10) and in the finite sample setting (see Fig.˜12). In both cases, we find that similar to SpecGD, Muon promotes learning different spectral components at a similar rate.
@@
We now show that this property of SpecGD to learn concepts at equal rate translates to superior generalization compared to GD in an imbalanced setting where least-significant spectral components of the moment matrices are associated to minority classes. Concretely, under data model (DM), the $k$ first eigenvectors of $\bm{\Sigma}_{{\bm{x}}{\bm{x}}}$ align with the class-mean vectors, ordered in decreasing class prior probability (see proof of Lemma 3.1). Intuitively, then, learning spectral components earlier during training should translate to generalization gains. Indeed, in Fig.˜2 (bottom), we observe that SpecGD attains lower minority-class and class-balanced loss compared to GD. The following theorem formalizes this intuition. We use continuous-time versions of GD and SpecGD here, i.e., gradient flow (GF) and spectral GF (SpecGF). The dynamics for GF are characterized by the approximation in Eq.˜2, while the SpecGF dynamics follow Proposition˜1, both with $\eta=1$ (see Sec.˜B.2 for details). We use $\bm{W}(t)$ and $\overline{\bm{W}}(t)$ to denote the continuous-time iterates and the corresponding singular value matrix, respectively.
@@
Next, we investigate whether these gains of SpecGD arise because it uses normalized updates while GD does not. To account for this difference, we compare SpecGD with NGD, the normalized version of GD. We first simulate the dynamics of NGD in the setting of Fig.˜2, and find that although NGD converges much faster than GD, its iterates $\overline{\bm{W}}_{t}$ follow the same trajectory as GD (Fig.˜2 (top)). Note that at the end of training, all three optimizers eventually converge to the same solution. However, importantly, we find that early on in training, SpecGD still outperforms NGD in terms of both minority-class and class-balanced loss (Fig.˜2 (bottom)). We formalize and prove this observation in the theorem below. See Sec.˜B.4 for the complete version and the proof.